# Linear-limit aging times of three monoalcohols

### v1

@@ Abstract
This paper presents data for the physical aging of the three monoalcohols 2-ethyl-1-butanol, 5-methyl-2-hexanol, and 1-phenyl-1-propanol. Aging is studied by monitoring how the dielectric loss at a fixed frequency in the kHz range equilibrates upon temperature jumps of a few Kelvin’s magnitude from equilibrium to equilibrium. The three alcohols differ in the Debye relaxation strength and how much the Debye relaxation time is different from that of the $\alpha$ process. We first demonstrate that single-parameter aging describes all data well and proceed to use this to identify linear-limit normalized aging relaxation functions. From the Laplace transform of these functions, the linear-limit aging loss-peak angular frequency defines the inverse of the linear aging relaxation time. This allows for a comparison to the temperature dependence of the Debye and $\alpha$ dielectric relaxation times of the three monoalcohols. We conclude that the aging response for 5-methyl-2-hexanol and 2-ethyl-1-butanol follows the $\alpha$ relaxation, not on the Debye process, while no definite conclusion can be reached for 1-phenyl-1-propanol because the Debye and $\alpha$ relaxations are here too merged to be distinguished.
@@ I Introduction
Physical aging is a characteristic feature taking place, in principle, in all glasses. It involves rearrangement of molecules below the glass transition temperature $T_{g}$, reflecting the fact that any glass continuously approaches the equilibrium liquid state [1, 2]. This can only be observed just below $T_{g}$, however, because aging is too slow at lower temperatures – aging of window glass is for instance not an issue. Understanding physical aging to the extent of being able to predict it quantitatively is important for applications of oxide glasses, polymers, metallic glasses, etc., both in production and subsequent use [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
The physical aging induced by a continuously changing temperature, e.g., as experienced during the production of glasses, is more difficult to model quantitatively than aging following a temperature jump. An ideal temperature jump starts from the glass-forming liquid annealed to equilibrium at the initial temperature, $T_{i}$, and changes the temperature rapidly to the final (annealing) temperature, $T_{a}$, which is kept constant for enough time that the system eventually reaches equilibrium at (we denote the final temperature by $T_{a}$ and not by $T_{f}$, which in aging contexts symbolizes the fictive temperature [16, 7]). The temperature change should be fast enough that virtually no relaxation takes place before the new temperature by heat conduction is established uniformly throughout the sample. After the jump, aging is monitored by measuring how some physical property, $X=X(t)$, approaches its equilibrium value. In the data reported below the property $X$ is the dielectric loss at a fixed frequency in the kHz range, but $X$ can also be, e.g., the density, the enthalpy, the dc conductivity, an elastic modulus, linear or nonlinear dielectric properties, etc [17, 3, 18, 5, 19, 20, 21, 22, 23, 24, 25, 26, 27]; aging of the structural relaxation time itself can be determined from strain-rate switching experiments, as recently convincingly demonstrated by Ediger and coworkers [28]. The only requirement is that an accurate measurement of $X$ can be accomplished within a short time window during which the system only changes its properties insignificantly.
@@
Achieving close-to-ideal temperature-induced aging conditions is not trivial, primarily because heat conduction is notoriously slow. We obtain such temperature jumps by working with a thin sample and by quickly changing the temperature utilizing a Peltier element close to the sample [29]. Recent alternative approaches to studies of physical aging include that of Henot et al., who used Ohmic heating to obtain extremely fast heating rates [30], and that of Cangialosi and co-workers utilizing the small sample size in fast scanning calorimetry to perform almost instantaneous temperature jumps [31, 32].
@@
Physical aging is highly nonlinear in the sense that different temperature jumps generally result in quite different normalized relaxation functions $R(t)$ defined by
@@
Here $X_{i}$ and $X_{a}$ are the equilibrium values of the monitored property at the initial temperature $T_{i}$ and the final (annealing) temperature $T_{a}$, respectively. It is henceforth assumed the jump takes place at $t=0$. Strong nonlinearity is usually observed even for temperature jumps of just 1% magnitude in the sense that down jumps are much faster and more stretched than up jumps of same magnitude to the same final temperature. This is referred to as the “asymmetry of approach”, with up jumps being “autocatalyzed” and down jumps “autoretarded” [17, 7, 33, 10, 12].
@@
The asymmetry of approach has been understood since long as an effect of the so-called material time, which controls the aging and is characterized by a rate of change that itself ages. The rate of aging of a temperature down jump decreases gradually and converges to the equilibrium aging rate at $T_{a}$, whereas for an up jump the aging rate increases from a low value to also eventually approach the $T_{a}$ equilibrium aging rate. A quantitative formalism has existed since 1971 [3], the so-called Tool-Narayanaswamy (TN) formalism, which accounts well for the nonlinearity of temperature jumps and, in fact, of any temperature protocol. An excellent account of the TN formalism was given by Scherer in his 1986 textbook [7].
@@
We used above the term “equilibrium aging rate”, which may seem like a contradiction in terms. Like any other response property, physical aging has a linear limit. Thus standard linear-response theory characterized by linearity and time-translational invariance applies for sufficiently small temperature jumps. This limit, which is approached only for temperature changes smaller than about 0.1%, corresponds to the situation in which the material time is a linear function of time throughout the aging process. A fluctuation-dissipation theorem exists also for linear physical aging, meaning that the aging response for very small jumps is determined by the equilibrium fluctuations [34]. It is in this sense an “equilibrium aging rate” may be defined [35].
@@
The linear limit of physical aging was reached experimentally by Riechers et al. [36], who reported data for aging following temperature jumps of varying amplitude, the smallest being just 10 mK [36]. That paper applied a simplified version of the TN formalism, the so-called single-parameter-aging (SPA) formalism from 2015 [37], to predict the results of nonlinear temperature jumps from those of small, linear jumps. Following Ref. 35, we use in this paper SPA to do the opposite: from nonlinear temperature jump data SPA is utilized to extract the difficult-to-measure linear aging relaxation function.
@@
The investigation presented below involves three monoalcohols. Monoalcohols are generally good glass formers, but differ from other organic glass-forming liquids by usually having a Debye process as the slowest and most intense dielectric relaxation process [38]. This is believed to reflect the dynamics of hydrogen-bonded structures, a process that is separate from the ordinary $\alpha$ relaxation associated with structural relaxation [38]. Most other linear-response functions have no – or merely very weak – relaxations at the dielectric Debye loss-peak frequency; they instead relax on more or less the time scale of the “ordinary” $\alpha$ relaxation process [38]. In particular, calorimetric measurements of monoalcohols generally correlate with the $\alpha$ relaxation rather than with the Debye process [39, 40].
@@
The question we address below is how physical aging proceeds in monoalcohols: Does aging take place on the time scale of the $\alpha$ process, as one may expect from the general identification of aging as reflecting structural relaxation [7, 12], or does aging proceed on the time scale of the slower dielectric Debye process? Since $\alpha$ linear responses may differ in their corresponding relaxation time but otherwise have identical temperature dependence [41, 42, 43, 44], we focus on comparing the aging relaxation time’s temperature dependence to those of the Debye and $\alpha$ relaxation times. In order to do this, however, it is necessary to have linear-response aging data at disposal, which as mentioned may be obtained from nonlinear aging data utilizing SPA.
@@
The investigation proceeds in two steps. First, we validate that the aging data conform to the SPA framework. This has previously been done for glycerol and for several van der Waals bonded liquids, as well as in computer simulations [37, 45, 46, 35, 47, 36, 48, 49], but not for monoalcohols that as mentioned have quite distinct properties. Once SPA has been validated, we proceed to use it to address whether physical aging as regards the temperature-dependence of the relaxation time follows the Debye process or the $\alpha$ process. Since the purpose is to compare to the temperature dependence of the dielectric Debye and the $\alpha$ relaxation times defined by $\tau=1/(2\pi\nu_{max})$ in which $\nu_{max}$ is the corresponding loss-peak frequency, we use the same equation to determine the “equilibrium” aging relaxation time $\tau_{R}$ from the loss-peak frequency $\nu^{R}_{max}$ of the frequency response obtained by Laplace transforming the SPA-extracted linear-limit aging relaxation function.
@@ II.1 Experimental details
We monitored aging after temperature jumps by continuously measuring the dielectric loss by means of a high-resolution Andeen-Hagerling AH2700A bridge. This was done at 1 kHz for 5M2H and at 10 kHz for 2E1B and 1P1P. The dielectric loss is the monitored property $X(t)$ in Eq. (1). Alternatively, one can use for $X(t)$ the real part of the dielectric constant; this gives more noisy data but leads to virtually the same linear aging relaxation times (data not shown). Our aging setup, which allows for accurate temperature control and rapid temperature jumps using a Peltier element and a thin liquid film (50 microns), is described in detail in Ref. 29.
@@
The TN formalism assumes the existence of a material time, $\xi$, with the property that all temperature jumps lead to the same normalized relaxation function $R(t)$ when given in terms of $\xi$. Thus according to TN, if one writes for any specific temperature jump $R(t)=\phi(\xi(t))$, the function $\phi$ is the same for all jumps whereas the functional form of $\xi(t)$ depends on the jump (usually strongly). If the jump goes from $T_{i}$ to $T_{a}=T_{i}+\Delta T$ where $\Delta T$ is the temperature change, SPA makes the following ansatz for the aging rate $\gamma$ [37, 50]
@@
Here $\gamma_{a}$ is the equilibrium linear aging rate at the final temperature of the jump, $T_{a}$, and $c$ is a material-specific constant. Note that the notation used follows that of Ref. 50 and differs from that of Ref. 37, which followed the old tradition in the aging field of defining $\Delta T$ as $T_{i}-T_{a}$, not as $T_{a}-T_{i}$, and which used a dimensionless version of the material-specific “non-linearity parameter” $c$. Note also that the material time – and thereby the aging rate $\gamma$ – is only defined within an arbitrary multiplicative constant.
Refs. 37, 51 showed how to use SPA to calculate the normalized relaxation function of one nonlinear temperature jump from that of another, Ref. 35 showed how SPA allows for calculating the linear limit of a nonlinear jump, and Ref. 36 showed how SPA can be used for calculating nonlinear jumps from linear ones. Although the reasoning is the same in all cases, we now for completeness summarize how SPA is used for calculating the linear-limit normalized aging relaxation function $R_{\rm lin}(t)$ from the normalized relaxation function $R(t)$ of a general, nonlinear jump.
For $R_{\rm lin}(t)$ the expression $R(t)=\phi(\xi(t))$ reduces to $R_{\rm lin}(t_{0})=\phi(\gamma_{a}t_{0})$ because $\xi=\gamma_{a}t_{0}$ in equilibrium at $T_{a}$ as well as very close to $T_{a}$ (the laboratory time of a linear-limit temperature jump is denoted by $t_{0}$ to distinguish it from the laboratory time $t$ of general, nonlinear jumps). Taking the time derivatives leads to $dR/dt=\phi^{\prime}(\xi(t))\gamma(t)$ and $dR_{\rm lin}/dt_{0}=\phi^{\prime}(\xi=\gamma_{a}t_{0})\gamma_{a}$, respectively. The ratio of these two equations evaluated at times $t$ and $t_{0}$ corresponding to the same value of $R$ (and thus the same material time $\xi$) is by Eq. (2) equal to $\exp(-c\Delta TR)$. For the ratio between the change of $t$ and $t_{0}$ corresponding to the same changes of $R$, $dR$, one thus has
$\frac{dt_{0}}{dt}\,=\,\frac{dR/dt}{dR/dt_{0}}\,=\,\frac{\gamma(t)}{\gamma_{a}}\,=\,\,e^{-c\,\Delta TR}\,,$ (3)
which means that
$dt_{0}\,=\,e^{-b\,\Delta TR}\,dt\,.$ (4)
For a temperature up jump one has $\Delta T>0$ and the time increment $dt$ needed for a given change of the normalized relaxation function, $dR$, is larger than the time increment $dt_{0}$ needed for the same change of the linear relaxation function at $T_{a}$. The opposite applies for a down jump. In both cases Eq. (4) provides the SPA recipe for calculating $R_{\rm lin}(t)$ from nonlinear jump data in the form of $R(t)$. This is done in a step-by-step manner starting at $t=0$ [37, 46, 36]. Since $R_{\rm lin}(t)$ for all jumps to $T_{a}$ must be the same function, the material-characteristic nonlinearity parameter $c$ can be determined by optimizing the collapse of the $R_{\rm lin}(t)$ functions calculated from different temperature jumps.

### v2

@@ Abstract
This paper presents data for the physical aging of the three monoalcohols 2-ethyl-1-butanol, 5-methyl-2-hexanol, and 1-phenyl-1-propanol. Aging is studied by monitoring the dielectric loss at a fixed frequency in the kHz range following temperature jumps of a few Kelvin’s magnitude, starting from states of equilibrium. The three alcohols differ in the Debye relaxation strength and how much the Debye process is separated from the $\alpha$ process. We first demonstrate that single-parameter aging describes all data well and proceed to utilize this fact to identify the linear-limit normalized aging relaxation functions. From the Laplace transform of these functions, the linear-limit aging loss-peak angular frequency defines the inverse of the linear aging relaxation time. This allows for a comparison to the temperature dependence of the Debye and $\alpha$ dielectric relaxation times of the three monoalcohols. We conclude that the aging response for 5-methyl-2-hexanol and 2-ethyl-1-butanol follows the $\alpha$ relaxation, not the Debye process, while no firm conclusion can be reached for 1-phenyl-1-propanol because its Debye and $\alpha$ processes are too close to be reliably distinguished.
@@ I Introduction
Physical aging is a characteristic feature of, in principle, all glasses. It involves the rearrangement of molecules below the glass transition temperature $T_{g}$, reflecting the fact that any glass continuously approaches the equilibrium liquid state [1, 2]. This can only be observed just below $T_{g}$, however, because aging is too slow at lower temperatures – aging of window glass is for instance not an issue. Understanding physical aging to the extent of being able to predict it quantitatively is important for applications of oxide glasses, polymers, metallic glasses, etc., both in production and subsequent use [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
The physical aging induced by a continuously changing temperature, e.g., as experienced during the production of glasses, is more difficult to model quantitatively than aging following a temperature jump. An ideal temperature jump starts from the glass-forming liquid annealed to equilibrium at the initial temperature, $T_{i}$, implementing a rapid change of temperature to the final (annealing) temperature, $T_{a}$, which is kept constant for enough time that the system eventually reaches equilibrium at (we denote the final temperature by $T_{a}$ and not by $T_{f}$, which in aging contexts symbolizes the fictive temperature [16, 7]). The temperature change should be fast enough that virtually no relaxation takes place before the new temperature by heat conduction is established uniformly throughout the sample. After the jump, aging is monitored by measuring how some physical property, $X=X(t)$, approaches its equilibrium value. In the data reported below the property $X$ is the dielectric loss at a fixed frequency in the kHz range, but $X$ can also be, e.g., the density, the enthalpy, the dc conductivity, an elastic modulus, a linear or nonlinear dielectric property, etc [17, 3, 18, 5, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Aging of the structural relaxation time itself can be determined from strain-rate switching experiments, as recently convincingly demonstrated by Ediger and coworkers [28]. The only requirement is that accurate measurement of $X$ can be carried out within a short time window during which the system only changes its properties insignificantly.
@@
Achieving close-to-ideal temperature-induced aging conditions is not straightforward, primarily because heat conduction is notoriously slow. We obtain such temperature jumps by working with a thin sample and by quickly changing the temperature by means of a Peltier element placed close to the sample [29]. Recent alternative approaches to studies of physical aging include that of Henot et al., who used Ohmic heating to obtain extremely fast heating rates [30], and that of Cangialosi and co-workers utilizing the small sample size in fast scanning calorimetry to perform almost instantaneous temperature jumps [31, 32].
@@
Physical aging is highly nonlinear in the sense that different temperature jumps generally result in quite different normalized relaxation functions, $R(t)$, defined by
@@
Here $X_{i}$ and $X_{a}$ are the equilibrium values of the monitored property at the initial temperature $T_{i}$ and the final temperature $T_{a}$, respectively. It is henceforth assumed the jump takes place at $t=0$. Strong nonlinearity is usually observed even for temperature jumps of just 1% magnitude in the sense that down jumps are much faster and more stretched than up jumps of same magnitude to the same final temperature. This is referred to as the “asymmetry of approach”, with up jumps being “autocatalyzed” and down jumps “autoretarded” [17, 7, 33, 10, 12].
@@
The asymmetry of approach has been understood since long as an effect of the so-called material time, which controls the aging and is characterized by a rate of change that itself ages. The rate of aging for a down jump in temperature decreases gradually and converges to the equilibrium aging rate at $T_{a}$, whereas for an up jump the aging rate increases from a low value to eventually also approach the $T_{a}$ equilibrium aging rate. A quantitative formalism has existed since 1971 [3], the so-called Tool-Narayanaswamy (TN) formalism, which accounts well for not only the nonlinearity of temperature jumps, but works for any temperature protocol. An excellent account of the TN formalism was given by Scherer in his 1986 textbook [7].
@@
We used above the term “equilibrium aging rate”, which may seem like a contradiction in terms. Like any other response property, however, physical aging has a linear limit. Thus standard linear-response theory characterized by linearity and time-translational invariance applies for sufficiently small temperature jumps. This limit, which is approached only for temperature changes smaller than about 0.1%, corresponds to the situation in which the material time is proportional to time throughout the aging process. A fluctuation-dissipation theorem exists also for linear physical aging, meaning that the aging response for very small jumps is determined by the equilibrium fluctuations [34]. It is in this sense an “equilibrium aging rate” may be defined [35].
@@
The linear limit of physical aging was recently reached experimentally by Riechers et al. [36], who reported data for aging following temperature jumps of varying amplitude, the smallest being just 10 mK [36]. That paper applied a simplified version of the TN formalism, the so-called single-parameter-aging (SPA) formalism from 2015 [37], to predict the results of nonlinear temperature jumps from data for small, linear jumps. Following Ref. 35, we use in this paper SPA to do the opposite: from nonlinear temperature jump data, SPA is utilized to extract the difficult-to-measure linear aging relaxation function.
@@
The investigation presented below involves three monoalcohols. Monoalcohols are generally good glass formers, but they differ from other organic glass-forming liquids by having a Debye process as the slowest and most intense dielectric relaxation process [38]. This is believed to reflect the dynamics of hydrogen-bonded structures, a process that is separate from the ordinary $\alpha$ relaxation associated with structural relaxation [38]. Most other linear-response functions have no – or only very weak – relaxations at the dielectric Debye loss-peak frequency; they instead relax on more or less the time scale of the “ordinary” $\alpha$ relaxation process [38]. In particular, calorimetric measurements of monoalcohols generally correlate with the $\alpha$ relaxation rather than with the Debye process [39, 40].
Since this paper studies the role of the Debye process in aging, the three monoalcohols were selected in order to have different degrees of separation between the Debye and the $\alpha$ relaxation processes as well as different magnitudes of their relative loss magnitudes. In all three cases the OH group is terminal or close to the end of the carbon chain. Two of the alcohols studied are regular simple alcohols, while the third one involves a phenyl group. For further studies it would be interesting to study a system of fixed-length monoalcohols for which the OH group location is varied systematically along the chain.
@@
The question we address below is how physical aging proceeds in monoalcohols: Does aging take place on the time scale of the $\alpha$ process, as one may expect from the general identification of aging as a consequence of structural relaxation [7, 12], or does aging proceed on the time scale of the slower dielectric Debye process? Since $\alpha$ linear responses may differ in their corresponding relaxation time but otherwise have identical temperature dependence [41, 42, 43, 44], we focus on comparing the aging relaxation time’s temperature dependence to those of the Debye and $\alpha$ relaxation times. In order to do this, however, it is necessary to have linear-response aging data at disposal, which are obtained from nonlinear aging data utilizing SPA.
The present work is not the first to address the possible role of the Debye process in physical aging. Previously, the focus was more on the absolute time scales than on their temperature dependence, however, and these studies did not attempt to identify the linear-limit aging time. Already in 2010 Gainaru et al. concluded that the dielectric aging of 2-ethyl-1-hexanol follows the $\alpha$ process [45]; in contrast, in the non-hydrogen-bonding liquid tri-butyl-phosphate (TBP) with an alcohol-like spectrum created by dielectric cross-correlation, the linear aging time follows the Debye-like mode associated with dielectric cross-correlations [46, 47]. Reference 48 studied mixtures of 2-ethyl-1-hexanol (2E1H) with 2-ethyl-1-hexyl bromide (2E1Br). For an equimolar mixture the Debye and $\alpha$ processes are separated by no less than four decades in frequency, and from a comparison of the involved time scales the authors concluded that physical aging is governed by the $\alpha$ process, not by the Debye process (see also Ref. 49). This is confirmed by the present study on pristine alcohols.
@@
Our investigation proceeds in two steps. First, we validate that the aging data conform to the SPA framework. This has previously been done for glycerol and for several van der Waals bonded liquids, as well as in computer simulations [37, 50, 51, 35, 52, 36, 53, 54], but not for monoalcohols that as mentioned have quite distinct properties making it highly nontrivial that they should comply to a simple SPA scenario. Once SPA has been validated, we proceed to use it to address whether physical aging in regard to the temperature dependence of the linear-limit aging time follows the Debye process or the $\alpha$ process. Since the purpose is to compare to the temperature dependence of the dielectric Debye and $\alpha$ relaxation times defined by $\tau=1/(2\pi\nu_{max})$ in which $\nu_{max}$ is the corresponding loss-peak frequency, we also use this equation to determine the “equilibrium” aging relaxation time $\tau_{R}$ from the loss-peak frequency $\nu^{R}_{max}$ of the frequency response (obtained by Laplace transforming the SPA-extracted linear-limit aging relaxation function).
@@ II.1 Experimental details
We monitored aging after temperature jumps by continuously measuring the dielectric loss by means of a high-resolution Andeen-Hagerling AH2700A bridge, which was done at 1 kHz for 5M2H and at 10 kHz for 2E1B and 1P1P. Thus the dielectric loss is the quantity $X(t)$ in Eq. (1). Alternatively, one can use for $X(t)$ the real part of the dielectric constant; this gives somewhat more noisy data but leads otherwise to virtually the same linear aging relaxation times (data not shown). Our aging setup, which allows for accurate temperature control and rapid temperature jumps using a Peltier element for thin samples (50 microns), is described in detail in Ref. 29.
@@
The TN formalism assumes the existence of a material time, $\xi$, with the property that all temperature jumps lead to the same normalized relaxation function $R(t)$ when this function is given in terms of $\xi$. Thus according to TN, if one writes for any specific temperature jump $R(t)=\phi(\xi(t))$, the function $\phi$ is the same for all jumps whereas the functional form of $\xi(t)$ depends (usually strongly) on the jump. If the jump goes from $T_{i}$ to $T_{a}=T_{i}+\Delta T$ where $\Delta T$ is the temperature change, SPA makes the following ansatz for the aging rate $\gamma$ [37, 55]
@@
Here $\gamma_{a}$ is the equilibrium linear aging rate at the final temperature of the jump, $T_{a}$, and $c$ is a material-specific constant. The notation used here is that of Ref. 55, which differs from Ref. 37 that followed the old tradition in the aging field of defining $\Delta T$ as $T_{i}-T_{a}$, not as $T_{a}-T_{i}$, and which used a dimensionless version of the “non-linearity parameter” $c$. Note also that the material time – and thereby the aging rate $\gamma$ – is only defined within an arbitrary multiplicative constant.
References 37, 56 showed how to use SPA to calculate the normalized relaxation function of one nonlinear temperature jump from that of another, Ref. 35 showed how SPA allows for calculating the linear limit of a nonlinear jump, and Ref. 36 showed how SPA can be used for predicting nonlinear jumps from linear ones. Although the reasoning is the same in all cases, we now for completeness briefly summarize how SPA is used for calculating the linear-limit normalized aging relaxation function, $R_{\rm lin}(t)$, from the normalized relaxation function $R(t)$ of a general, nonlinear jump.
For $R_{\rm lin}(t)$ the expression $R(t)=\phi(\xi(t))$ reduces to $R_{\rm lin}(t_{0})=\phi(\gamma_{a}t_{0})$ because $\xi=\gamma_{a}t_{0}$ in equilibrium at $T_{a}$ as well as very close to $T_{a}$; here and henceforth the laboratory time of a linear-limit temperature jump is denoted by $t_{0}$ to distinguish it from the laboratory time $t$ of general, nonlinear jumps. Taking the time derivatives leads to $dR/dt=\phi^{\prime}(\xi(t))\gamma(t)$ and $dR_{\rm lin}/dt_{0}=\phi^{\prime}(\xi=\gamma_{a}t_{0})\gamma_{a}$, respectively. The ratio of these two equations evaluated at times $t$ and $t_{0}$ corresponding to the same value of $R$ (and thus the same material time $\xi$) is by Eq. (2) equal to $\exp(-c\Delta TR)$. For the ratio between the time increments $dt$ and $dt_{0}$ corresponding to the same decrease of $R$, $dR$, one thus has
$\frac{dt_{0}}{dt}\,=\,\frac{dR/dt}{dR/dt_{0}}\,=\,\frac{\gamma(t)}{\gamma_{a}}\,=\,\,e^{-c\,\Delta TR}\,.$ (3)
This implies that
$dt\,=\,e^{c\,\Delta TR}\,dt_{0}\,.$ (4)
For a temperature up jump one has $\Delta T>0$ and the time increment $dt$ needed for a given change of the normalized relaxation function, $dR$, is larger than the time increment $dt_{0}$ needed for the same change of the linear relaxation function at $T_{a}$. The opposite applies for a down jump. In both cases Eq. (4) provides the SPA recipe for calculating $R_{\rm lin}(t)$ from nonlinear jump data in the form of $R(t)$. This is done in a step-by-step manner starting at $t=0$ by transforming the time axis of $R(t)$ to that of $R_{\rm lin}(t)$ by means of Eq. (4) [37, 51, 36]. Since $R_{\rm lin}(t)$ for all jumps to $T_{a}$ must be the same function, the material-characteristic nonlinearity parameter $c$ can be determined by optimizing the collapse of the $R_{\rm lin}(t)$ functions for jumps to the same temperature. This is done for all final temperatures $T_{a}$; $c$ is then the nonlinearity parameter that results in the best overall collapse.